UT 20170115 LUCI1 AO Nighttime Commissioning

Observer LUCI-Team: JHeidt, DThompson (Tuson)
Observer AO: GTayler, DMiller
Telescope Operator: GBechetti

Plan for the night

- We are clouded out thus we will do closed-dome-tests

- Subarray readout-tests GEIRS

- AFC-ima LUCI1

Summary

- Subarray-tests by Dave successfull

- Not much AFC-testing, had a problem with the compensation mirror, which Alexander solved.

Details

All times are UT
Files are luci1.20170115.NNNN.fits

00:35 Sunset, Weather: overcast, T=0 deg, Hum=100%, Wind=15-20 m/s

00:55 Dave is starting GEIRS-tests

01:15 Dave is done with his tests

01:20 Geno sent a test preset which failed, needs to evaluate, restartet PCS et al.,

02:00 Start with AFC-tests, problems with compensation mirror, call TSargent

03:00 Tom wasn't able to solve problem, need to wait until Alexander is available

07:30 Alexander checked the issue and solved it, from his E-Mail:

We have prepared the flexure compensation code for telescope elevation
more 90 degree in December but it was not deployed.

Now the telescope has elevation like 90.0004. We are out of lookup table ranges.

I have deployed update:

if (currentElevation > 90) currentElevation = 90.0;
if (currentElevation < 0) currentElevation = 0.0;

@Tom, also all your changes from last days were deployed.

10:45 We did some tests afterwards, everything is fine now.

-- %USERSIG{JochenHeidt - 2017-01-15}%
Topic revision: r4 - 16 Jan 2017, JochenHeidt
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback